OpenStudy (ccastudentgrade10):

help pleasee!!! simplify the expression? -5+i/2i please explain i have a bunch of these to do!

OpenStudy (ccastudentgrade10):

@Nishant_Garg @xMissAlyCatx ??

OpenStudy (dayakar):

\[\frac{ -5+i }{ 2i }\] is this ur question

OpenStudy (ccastudentgrade10):

yeah

OpenStudy (dayakar):

is the division sigh includes 5 or not clarify

OpenStudy (ccastudentgrade10):

what?

OpenStudy (dayakar):

is it \[5 + \frac{ i }{ 2i }\]

OpenStudy (ccastudentgrade10):

no its the way you put it before

OpenStudy (xmissalycatx):

Combine like terms. 1/2 + 5i/2

OpenStudy (ccastudentgrade10):

so .5 + 2.5i?

OpenStudy (dayakar):

\[\frac{ -5 }{ 2i }+\frac{ i }{ 2i }\]

OpenStudy (xmissalycatx):

@Hero

OpenStudy (xmissalycatx):

@Hero , can you please help me explain this?

OpenStudy (dayakar):

\[=2.5i +0.5\]

OpenStudy (ccastudentgrade10):

is that the whole answer? im confused of the steps.

OpenStudy (dayakar):

-2.5i+0.5

OpenStudy (dayakar):

5/2 =2.5

OpenStudy (ccastudentgrade10):

but how are we dividing it by 2?

OpenStudy (xmissalycatx):

What you want to do is multiply EVERYTHING by -i, because if it is a negative, it will cancel out on the bottom.

OpenStudy (ccastudentgrade10):

so -5 + i/2i would be 5i/-2i?

OpenStudy (dayakar):

no

OpenStudy (ccastudentgrade10):

im lost can someone explain it like all in one message?

OpenStudy (xmissalycatx):

Let me draw it out.. hold on.

OpenStudy (dayakar):

\[\frac{ x+y }{ 2 }= \frac{ x }{ 2 }+\frac{ y }{ 2 }\]

OpenStudy (ccastudentgrade10):

but the bottom is 2i

OpenStudy (xmissalycatx):